Elle Inc. has two types of handbags: standard and custom. The controller has decided to use a plant-wide overhead rate based on direct labour hours. The president has heard of activity-based costing and wants to see how the results would differ if this system were used. Two activity cost pools were developed: machining and machine set-up. Presented below is information related to the company's operations.

Standard

Custom

Direct labour costs

$50,000

$100,000

Machine hours

500

1,000

Set-up hours

100

400

Total estimated overhead costs are $277,500. The overhead cost allocated to the machining activity cost pool is $177,000, and $100,500 is allocated to machine set-up activity cost pool.

Please answering the following questions:


(a) Calculate the overhead rate using the traditional approach.

Answer% of direct labour cost.


(b) Calculate the overhead rates using the activity-based costing approach.

Machining: $________ per machine hour
Machine setup: $________per setup hour.

(c) Calculate the allocation of overhead to the standard handbag under both approaches:

Traditional: $_________
Activity-Based: $________

(d) Calculate the allocation of overhead to the custom handbag under both approaches:

Traditional: $________
Activity-Based: $________
Answer #1
Estimated overhead = Predetermined overhead rate
Direct labor costs
277500/($50,000 + $100,500) = 200% of direct labor cost
Activity cost pools Cost drivers Estimated overhead
Machining Machine hours 177000
Machine setup Set up hours 100500
Activity-based overhead rates
Machining: Machine setup:
177000/(500+1000) = $118 per machine hour 100500/500 = $201 per setup hour
Traditional costing Standard Custom
$50,000 X 200% 92193
$100,000 X 200% 184385
92193 184385
Activity-based costing Standard Custom
Machining:
500 X $118 59000
1,000 X $118 118000
Machine setup:
100 X $201 20100
400 X $201 80400
Total 79100 198400